OpenStudy (abdumu16):

Meg's pension plan is an annuity with a guaranteed return of 6% per year (compounded quarterly). She would like to retire with a pension of $10,000 per quarter for 25 years. If she works 45 years before retiring, how much money must she and her employer deposit each quarter?

OpenStudy (coacoapuffprincess):

Ok

OpenStudy (coacoapuffprincess):

One second

OpenStudy (coacoapuffprincess):

she needs 50,000*(1-1.015^-100)/0.015 at the time of retirement which should = FV of deposits at the time of retirement = PMT(1.015^164-1)/0.015 equating the two, PMT = $3689.96 <

OpenStudy (coacoapuffprincess):

Meg's pension plan is an annuity with a guaranteed return of 5% per year (compounded quarterly). She would like to retire with a pension of $10,000 per quarter for 15 years. If she works 35 years before retiring, how much money must she and her employer deposit each quarter? HINT [See Example 5.] (Round your answer to the nearest cent.)

OpenStudy (coacoapuffprincess):

U should shoot for the first one though

OpenStudy (abdumu16):

Thank you @coacoapuffprincess But webassign shows that the answers is wrong I have no idea how to do this one

OpenStudy (abdumu16):

PV = (1-(1+i)^-n )/i PV = FV PMT = FV i/((1+i)^n)-1 I solved it with this equation @cocoapuffprincess